Please help with 4-8Determine the amount of sales (units) that would be necessary under Break-Even Sales Under Present and Proposed Conditions Da4. Compute the break-even sales (units) under the proposed program for the following year. Enter the final answers rounded to
Answer #1

4.

Break even sales under proposed program = $1,922,400+52,800 / $16.59 = 119,060 units

5.

Desired sales units = $1,975,200+71,200 / $16.59 = 123,351 units

6.

Maximum income from operations :

Increase in sales units = $528,000 / $48 = 11,000 units

Sales (120,150+11,000)*$48 $6,295,200
Variable cost (131,150*$31.41) 4,119,422
Contribution margin 2,175,778
Fixed costs 1,975,200
Income from operations $200,578

7.

Income or loss from operations at current sales level = $71,200 - 52,800 = $18,400

8. In favor of the proposal because of the possibility of increasing income from operations.